The value of ∆G•for phosphorylation  of the glucose in the glycolysis is 13.8 Kilo joule per mole find the value of kc  at 298 Kelvin

Grade:11

1 Answers

Arun
5 years ago
ΔG=-2.303RT log Kc
where R =8.314jk^-1mol^-1;T=298k;
Given
ΔG=13.8KJ/Mol but has value of R is in joules,we take it as 13.8×10³j/mol
therefore 13.8×10³=-2.303×8.314×298logKc
log Kc =-2.4185
antilog=0.2661×10^-2